OpenStudy (anonymous):

Cell phones are becoming very popular at Poplar Middle School. In 2000, there were 15 students with cell phones. Each year the number of students with cell phones has increased by 20%. At this rate, how many students will have cell phones in 2007? a. 44 students b. 54 students c. 35 students d. 38 students

OpenStudy (anonymous):

@ladycomical

OpenStudy (anonymous):

@bettni

OpenStudy (anonymous):

@litchlani

OpenStudy (anonymous):

@Diwakar

OpenStudy (anonymous):

It appears to be a problem of geometric progression. Each year, no. of phones increases by 20%. If in nth year no of phones were xn. Then in the year x(n+1), no of phones will be x(n+1)=xn+(20% of xn) =1.2*xn Similarly x(n+2) =x(n+1) + (20% of x(n+1)) =1.2*x(n+1) = 1.44*xn So, the series will be : xn, 1.2 xn, 1.44 xn We got common ratio as 1.2 first term (year 2000)=15 We have to find the 8th term (year 2007) Use \(\large{a_{n} = a_{1} r^{n-1}}\) with a1=15 n=8 r=1.2

OpenStudy (anonymous):

I am getting the answer as ~ 54